2019-04-21

Network disturbance in Norrtälje

Privatpersoner hänvisas till sina tjänsteleverantörer för mer information
End customers are kindly referred to their individual service providers.

Timezone: CEST
Incident start: 2019-04-21 04:05
Incident description: Network disturbance
Time of repair: 2019-04-23 12:00

Affected Users: Private segment
Affected Services: Several services
Affected geographical area: Norrtälje, 76112, 76111, 76197, 76198, 76194, 76193, 76015

Symptom: Outage
Impact: Medium
Cause of disturbance: Suspected fiber damage
Current progress report:
2019-04-21 04:05 – Alarms raised, Troubleshooting started
2019-04-21 06:30 – Technician is dispatched to site for troubleshooting.
2019-04-21 09:00 – Technician going towards secondary site for further troubleshooting.
2019-04-21 10:30 – Technician has been able to troubleshoot secondary site, possible cause is hardware failure
2019-04-21 11:30 – Technician arrives at IP-Only NOC to receive a SFP replacement.
2019-04-21 11:40 – Technician on his way to secondary site to replace SFP.
2019-04-21 13:10 – Technician swapped SFP and cleaned, no change in status.
2019-04-21 13:40 – Fiberwash and port oversight done, no change.
2019-04-21 14:10 – Thirdline activated and NOC-Manager informed.
2019-04-21 15:00 – On-call fiber specialist advised.
2019-04-21 15:30 – Wavelength specialist called in for further troubleshooting.
2019-04-21 21:00 – Technician is dispatched to change circuit to another wavelength.
2019-04-21 03:30 – Technician has changed wavelength and SFP’s but the problem persists.
Troubleshooting has been ongoing during the night. A new technician needs to be sent out to continue troubleshooting.
2019-04-22 07:00 – A new technician has been allocated.
2019-04-22 09:30 – Technician is en route.
2019-04-22 11:30 – Technician has measured a suspected fiber damage 6981 meters from site. Third level team engaged to estimate the location of the damage.
2019-04-22 13:00 – Issue has been escalate to the responsible team for installation in the area. A new team is en route. 15 minutes ETA.
2019-04-22 14:30 – Search continues for location, the length of the cable is being inspected but is delayed due to difficult terrain.
2019-04-22 15:30 – Search continues for fault location. Two tech teams are troubleshooting onsite.
2019-04-22 16:30 – Search continues for fault location.
2019-04-22 17:30 – New plan to repair the link with most impacted customers, we will reroute to another path and continue looking for the fiber cut with less impact on end customers. ETR for reroute is set to 2019-04-22 19:30.
2019-04-22 19:30 – There is a small delay, our technician is on the last site and swapping the necessary patches/hardware right now. ETR 2019-04-22 20:30
2019-04-22 21:30 – Services are being restored. We will continue to troubleshoot the root cause to restore the switch with less affected end customers. Our plan now is to move that switch tomorrow. Next update will come at 2019-04-23 08:30

2019-04-23 08:30 – Troubleshooting continues
2019-04-23 10:00 – Troubleshooting ongoing.
2019-04-23 11:00 – Technicians have retrieved new hardware and is heading back to the site to continue the troubleshooting. ETA 30 minutes.
2019-04-23 12:00 – Hardware have been replaced and configuration done, all services restored.

Add our RSS to your e-mail or app: http://www.ip-only.se/feed/?cat=6

Privatpersoner hänvisas till sina tjänsteleverantörer för mer information
End customers are kindly referred to their individual service providers.

Till startsidan